OpenStudy (anonymous):

find in the R the solution set x^2-5x-6=0

OpenStudy (anonymous):

(x -6) (x + 1) = 0 x= 6 or -1. R = { x / x = 6 or x = -1} R = { x/ x \[\in\] {-1,6} }

OpenStudy (anonymous):

thnxxxxxx but hw did u get it ANA2 ?

OpenStudy (anonymous):

(x -6) (x + 1) = 0 If a*b = 0 either a or b are zero, or both. if (x-6) = 0 x = 6 if x + 1 = 0 x = -1.

OpenStudy (anonymous):

thnxxxx